Your portfolio contains $12,000 of BMO stock, which has a beta of 1.1, and $18,000 of Lululemon stock, which has a beta of 1.4. What is the beta of your portfolio?
Answer #1

Total value = 12,000 + 18,000 = $30,000

Beta = (12,000 / 30,000)*1.1 + (18,000 / 30,000)*1.4

Beta = 0.44 + 0.84

Beta = 1.28